DIRC Operations Report
Wednesday 2007/8/22 to Wednesday 2007/8/29
Oncall: Rolf Andreassen (Justine Serrano on backup)
Fast Monitoring: Rolf Andreassen, Justine Serrano
Luminosities this week
PEP peak lumi: ~1.17 x 1034
(last week: ~1.10 x 1034)
PEP delivery: 3.64 fb-1
(last week: 6.00 fb-1)
BaBar integrated lumi:3.47 fb-1
(last week: 5.68 fb-1)
Run 6 total delivery: 85.91 fb-1
Run 6 total BaBar recorded: 82.49 fb-1 total, including 7.89 fb-1 off-peak data
PEP activities/issues this week
BaBar activities/issues this week
08/27-29 : The narrower timing window for the SVT was introduced in an effort to reduce deadtime.
DIRC activities/issues this week
Some noisy channels, which makes the shifters a little nervous.
08/27: Jochen and Chris tested the new FEX code with three colliding beam runs. Fast Mon and Prompt Calibration plots looked ok.
08/29: Matt replaced the water filters.
See DIRC logbook for details
Lumi and Background this week
- Take a standalone run at the next 1-2 hour down time to investigate the Xmas trees.
- Next Tuesday Linac will shutdown at 6am and PEP will dump beams. A global calibration will be taken. Later on a ADC calibration and a random trigger run will be taken. Then HV will be turned off.
- All this will be done in October when the doors are open.
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
- Unplug christmas tree in sector 5 and sector 11.
- Replace air filters on Wienner crates.
- Investigate the crate of sector 8 which has again killed the first 8 channels of DFB 7.
| Equipment || # used || # spare (*) || Comments |
| DFB || 168 || 18
6 good (3F,4E,64,66,68,69),
3 potentially good (6E,57,9C)
9 bad (05,70,6C,78,40,88,95,87,5A) in France
| DCC || 12 || 7
3 good (12,31,5)
2 bad (25,9) in France
- Rolf will be oncall until mid-September, then Jose will take over.
- Justine has left, Loic will arrive in October.